OpenStudy (anonymous):

(3)^1/2y^2−4y−7(3)^1/2=0 use the discriminant to determine the number of solutions of the quadratic equation, and whether the solutions are real or complex. Note: It is not necessary to find the roots; just determine the number and types of solutions

OpenStudy (anonymous):

Assuming this to be:\[\sqrt{3}y^2-4y-7\sqrt{3}=0\] Then the discriminant is the bit in the square root of the quadratic formula. Here it's: \[4^2-(4*\sqrt{3}*-7\sqrt{3})\]= 16+84 = 100. If the disciminant is below 0, there are no real solutions; if it is equal to 0, there is a repeated root; since this is above 0, there are two real solutions.

OpenStudy (anonymous):

so is this real or complex?

OpenStudy (anonymous):

Quote: "since this is above 0, there are two real solutions"